OpenStudy (anonymous):

Is this correct? 3/x + 3/x+1 + 3/x+2. I got the CD x(x+1)(x+2) and eventually I got (3x+3)(3x+2)/x(x+1) + (3x^2+6)/x(x+1)(x+2) + (3x^2+3)/x(x+1)(x+2)

OpenStudy (anonymous):

This isn't my final answer i just want to know before i go on. \[\frac{ (3x+3)(3x+6) }{ x(x+1)(x+2) } + \frac{ 3x ^{2}+6 }{ x(x+1)(x+2) } + \frac{ 3x ^{2}+3 }{ x(x+1)(x+2) }\]

OpenStudy (anonymous):

its wrong,the first numerator should be 3(x+1)(x+2)=(3x+3)(x+2) the second 3x(x+2)=3x^2+6x and the third 3x(x+1)=3x^2+3x

OpenStudy (anonymous):

OK thanks :)

OpenStudy (anonymous):

So do I just combine the numerators next?

OpenStudy (anonymous):

yes and see if the resulting numerator is factorizable

OpenStudy (anonymous):

Do i multiply the (3x+3)and (x+2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got (x+1) (x+1)

OpenStudy (anonymous):

\[\frac{ (x+1) }{ x(x+2) }\] is this the final answer?

OpenStudy (anonymous):

3/(x) + 3/(x+1) + 3/(x+2) 3(x+1)(x+2)+3x(x+2)+3x(x+1) =------------------------------------ x(x+1)(x+2) (3x+3)(x+2)+3x^2+6x+3x^2+3x =---------------------------------- x(x+1)(x+2) 12x^2 +18x+6 =-------------------------------- x(x+1)(x+2) 6(2x^2+3x+1) =------------------------------- x(x+1)(x+2) 6(x+1)(2x+1) =---------------------- x(x+1)(x+2) 6(2x+1) = ---------------- x(x+2)